answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of drawing the six clubs from a standard deck of cards?

The answer depends on:

  • whether or not the cards are drawn at random,
  • whether or not the cards are replaced before drawing another,
  • how many cards are drawn.
If 45 cards are drawn, without replacement, the event is a certainty.

What is the probability that a human sperm cell Will carry a X chromosome?

The answer will vary from person to person. It will depend on the individual as well as his age.

Why can't probability be less than 0?

Because probability is measured from zero to one, where zero means something cannot happen and one meaning it will definitely happen. This can also be looked at as there being a 0% chance or 100% chance of it happening. So obviously there can not be less of a chance than it definitely not happening, so it cannot be less than zero.

A number is chosen 1 to 25 find the probabillity of selecting a number greater than 18?

Probability is ratio of the events you want over the total number of events. There are 7 numbers greater than 18 and you have 25 total options, so your probability is 7/25 or 28%.

What is the sample space for choosing the letter from the word probability?

The sample space consists of the letters of the word "PROBABILITY" = {P,R,O,B,A,I,L,T,Y}

What is the probability of selecting March?

The answer depends on what you are selecting from. If you are selecting months in which the equinoces occur, the probability is 0.5

What is a biased coin probability?

A biased probability is one where not every outcome has the same chance of occurring. A biased coin is one where one side, the "heads" or "tails" has a greater probability than the other of showing.

A coin which has a centre of gravity closer to the tails side than the heads side would be biased in that heads is more likely to show than tails.

The size of coin can have an effect on the probability of heads and tails - during the Royal Institute Christmas lectures in the 1990s demonstrating probability a large version of the pound coin was made to be able to allow the audience to see it being tossed - on the broadcast (and tape) version it landed and stayed on its edge! showing the probability of heads = tails ≠ ½; the probability of heads = probability of tails, but they are actually slightly less than ½ as the coin could land on its edge and stay there - with a standard size coin, if it lands on its edge it takes very little for the centre of gravity to shift outside the base of the edge and for the coin to fall over, but with a very large similar coin (ie one scaled up [proportionally] in lengths) it can take quite a bit before the centre of gravity goes outside the base if it lands on its edge which forces it to fall over (plus there will be a "significant" rise in the centre of gravity to do so, thus favouring stability on an edge which does not exist in the standard, small, sized version of the coin).

What is the purpose of finding standard deviation?

The standard deviation of a set of data is a measure of the random variability present in the data. Given any two sets of data it is extremely unlikely that their means will be exactly the same. The standard deviation is used to determine whether the difference between the means of the two data sets is something that could happen purely by chance (ie is reasonable) or not.

Also, if you wish to take samples of a population, then the inherent variability - as measured by the standard deviation - is a useful measure to help determine the optimum sample size.

A card is randomly drawn from an ordinary deck of playing cards what is the probability that the card is a king or a diamond?

You randomly select one card from a 52-card deck. Find the probability of selecting the king of diamonds or the jack of

The theoretical probability of a given event is 9 13 explain what each number represents?

The naive answer is that the random experiment (or trial) can be modelled theoretically. This model has 13 equally likely outcomes and that 9 of them are favourable to the event. That is, 9 out of 13 equally likely outcomes involve the event occurring.

A slightly more sophisticated explanation is that there are a multiple of 13 equally likely outcomes and the same multiple of 9 are favourable to the event under study. For example, the probability of randomly drawing a non-honour from a regular deck of cards. The total number of possible outcomes is 52 (= 13*4) and the number of favourable outcomes is 36 (= 9*4). The probability, 36/52 has then been simplified to 9/13.


This answer still assumes that each individual outcome is equally likely but that is not necessary. The experiment could involve trying to hitting one of two targets: a 2*2 and a 3*3 (in the same square units of length) targets. If anywhere in the two targets has the same probability, then the theoretical probability of hitting the bigger target is 9/(9+4) = 9/13.

How do you make predictions using experimental probability?

Examples like the propability for raining tommorrow will 1/2 may or may not happen probability is called possibility

What are the odds of rolling a double one in dice throwing?

Possible outcomes with one cube = 6

Possible outcomes with 2 cubes = (6 x 6) = 36

Probability of a double-one (snake eyes) = 1/36 = 2.78% (rounded)

Odds: 35 to 1 against

Find the number of distinguishable permutations of letters in the word appliance?

The distinguishable permutations are the total permutations divided by the product of the factorial of the count of each letter. So:

9!/(2!*2!*1*1*1*1*1) = 362880/4 = 90,720

How many ways can 6 different books be stacked on a shelf?

First book can go in any of 6 places, second in any of the 5 remaining and so on until fifth book has 2 options.

There are therefore 6 x 5 x 4 x 3 x 2 ie 720 possible arrangements.

A coin is tossed 100 times how many times would you expect to get heads?

The probability of a heads is 1/2. The expected value of independent events is the number of runs times the probability of the desired result. So:

100*(1/2) = 50 heads

Trending Questions
How do you calculate the third quartile for a mean of 110 and a standard deviation of 15? How many possible outcomes are there when flipping a coin 8 times? What is 1-2 of 1-3 cup? Is cluster sampling random? What is the probability of rolling a number less than 5 AND an odd number on one die? The probability of a single hazard occurring is? Your father is a twin can you have twins? What is the probability that the number on the ball is greater than or equal to 26? The probability that kelly goes to the school is 0.39 What is the probability that kelly will not go to school? In a bag containing 6 purple marbles 19 green marbles 5 blue marbles 19 yellow marbles and 17 red marbles what is the probability of pulling out a yellow marble? Hazard severity categories are assigned? What is the probability of flipping no heads when you flip three identical coins? question . If two normal dice are thrown together, what is the probability of getting a sum of 6? How do you arrange the word rirorm? What is the probability that a continuous random variable takes on a single value? What has a head a tail and brown? If there are 12 possible outcomes what is the lowest probability that will still make an outcome likely? Find the probability of rolling a sum of 12 with two dice? What is the probability of rolling a sum of 9 or less Suppose you roll two dice one time only? Boddy has a bag full of marbles that he keep in his room he has 35 red marbles and 25 green marblesfind the ratito of red marbles and put it in its simplest form?